Fermi National Laboratory


Accelerator Update

Wednesday December 10
The day shift began with the Tevatron (Tev) in access, stacking in the Antiproton Source (Pbar), and with beam to MiniBoone.

Studies were conducted in the Main Injector (MI) and in Pbar for much of the day and evening shift.

Switchyard 120 (SY120) studies began at 3:00 PM.

SY120 and MI studies ended at 11:11 PM.

Thursday December 11
The midnight shift began with Pbar stacking and conducting studies, and with MiniBoone taking beam.

The H- source failed at 3:54 AM.

At 6:33 AM they changed to the I- source and did some Linac tuning and Booster injection tuning.

Five Pbar shots were sent to the Recycler Ring (RR) at 11:00 AM.

Accesses into Linac, Booster, Pbar, and MI-13 occurred during the afternoon.

Conning Tower replacement was complete at 2:30 PM and leak checking is beginning.

All accesses were completed by 4:30 PM. Tev is still in access for leak checking.

At 5:00 PM Linac and Booster beam returned, tuning for Miniboone and SY120 began, and Pbar studies are taking place.

Sy120 studies ended at 11:58 PM.

Friday December 12
The midnight shift began with Pbar stacking and conducting studies, Booster studies, and with MiniBoone taking beam.

Worked on loss problems in Booster and MiniBoone most of the midnight shift.

Today's Plans
Continue with the Tev conning tower checkout.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/12/2003   email Fermilab